OpenStudy (summersnow8):

On a part-time job, you are asked to bring a cylindrical iron rod of density 7800kg/m3 , length 92.5cm and diameter 2.65cm from a storage room to a machinist. Calculate the weight of the rod, w. The acceleration due to gravity, g = 9.81m/s2 .

OpenStudy (anonymous):

"Machinist" reminds me of Christian Bale. :P By the way, this is a simple question, provided you know how to calculate the volume of a cylinder. Then, mass = density x volume. And weight = mass x g

OpenStudy (anonymous):

umm., you know the formula ? pie r square times length = vol of cylinder. And , yes be careful with the units.

OpenStudy (summersnow8):

i got a huge number....

OpenStudy (anonymous):

do you want to cross check ? give me the ans,

OpenStudy (summersnow8):

what do you mean by cross check...? so for V=pir^2L I have 5.102, then m=pv then i get 39794..... then multiply that by 9.81 i get 390379

OpenStudy (anonymous):

aw,. cross check means comparing answers. Well, mine is different. Wait, I'll show the steps.

OpenStudy (summersnow8):

ok

OpenStudy (anonymous):

\[7800\times \pi \times0.925\times0.01325^{2}\times9.81\] got it, you screwed up with the units !

OpenStudy (summersnow8):

wait what is the equation for volume?

OpenStudy (summersnow8):

why are you including density?

OpenStudy (anonymous):

\[Vol = \pi \times r ^{2}\times l\]

OpenStudy (anonymous):

mass = density x volume , without density wherefrom will you get the mass ? Weight requires mass ! And, mass is given indirectly here by density and volume Got it ?